跳到论文正文
W Papers独立论文阅读器 arXiv:2608.26996 论文正文 计算阅读时间 · 0%
引用这篇论文根据当前论文元数据生成,正式投稿前建议核对。

    
  
License: CC BY 4.0
arXiv:2608.26996v1 [math.PR] 27 Aug 2026

A note on shifting distributions via Poisson races

Amir Yehudayoff Address: Department of Computer Science, The University of Copenhagen and Department of Mathematics, Technion-IIT Email address: amir.yehudayoff@gmail.com
Abstract.

This expository note is about simulating a target distribution QQ from observations of a proposal distribution PP. In the model suggested by Harsha, Jain, McAllester and Radhakrishnan, we observe an infinite sequence of i.i.d. samples X1,X2,…X_{1},X_{2},\ldots from PP. The goal is to find some index I∈{1,2,…}I\in\{1,2,\ldots\} such that XI∼QX_{I}\sim Q while minimizing ⁡E​log⁡I\mathop{\mathbb{}}{E}\log I. Following the Poisson-race approach developed by Maddison, Li and El Gamal, and others, this note shows that if D(Q||P)<∞D(Q||P)<\infty then there is a PP to QQ simulator II such that E[logI]≤D(Q||P)+1.45∥Q−P∥1\mathop{\mathbb{}}{E}[\log I]\leq D(Q||P)+1.45\|Q-P\|_{1}. In the other direction, for every PP to QQ simulator II, the cost is at least E[logI]≥12max{D(Q||P),∥Q−P∥1}\mathop{\mathbb{}}{E}[\log I]\geq\frac{1}{2}\max\{D(Q||P),\|Q-P\|_{1}\}.

1. Introduction

The problem of simulating a target distribution QQ given access to some other distribution PP has been studied for a long time. It was studied by von Neumann [13] who introduced rejection sampling, and by Wyner [14] who introduced the common information problem. The specific form of the problem described below was first suggested, to the best of my knowledge, in the work of Harsha, Jain, McAllester and Radhakrishnan [6]. But this type of problem already implicitly appeared in the work of Hajek and Pursley [5], and a close variant of it was termed the functional representation lemma by El Gamal and Kim [3] and the strong functional representation lemma by Li and El Gamal [9]. This note is too brief to provide a comprehensive description of the history (for more details, see the references above and also [1, 4, 11] and references therein).

There are two distributions P,QP,Q over some finite set11 1 We focus on the case of a finite set for simplicity. 𝒳\mathcal{X}. We observe an infinite sequence of i.i.d. samples X1,X2,…X_{1},X_{2},\ldots from PP. Our goal is to find some index I∈{1,2,…}I\in\{1,2,\ldots\} such that XI∼QX_{I}\sim Q; that is, Pr[XI=x]=Q(x)\Pr[X_{I}=x]=Q(x) for every x∈𝒳x\in\mathcal{X}. We call such an index II a simulator (it can also be randomized).

Here is an important example: a two-player game with public shared randomness. Alice knows some distribution RR on pairs 𝒳×𝒴\mathcal{X}\times\mathcal{Y}, and Bob knows the marginal distribution R𝒴R_{\mathcal{Y}}. Alice gets as input X∼R𝒳X\sim R_{\mathcal{X}} and her goal is to communicate some information to Bob that will allow Bob to compute YY with the property that (X,Y)∼R(X,Y)\sim R. The two extreme cases are (1) when R=R𝒳×R𝒴R=R_{\mathcal{X}}\times R_{\mathcal{Y}} is a product distribution and (2) when 𝒳=𝒴\mathcal{X}=\mathcal{Y} and RR is supported on the diagonal {(x,x)}\{(x,x)\}. The players can use a simulator to solve this problem. The players interpret the shared randomness as infinitely many i.i.d. samples Y1,Y2,…Y_{1},Y_{2},\ldots from P:=R𝒴P:=R_{\mathcal{Y}}. Alice gets XX and constructs a simulator II for Q⁡(y):=R⁡(y|X)Q(y):=R(y|X). She sends II to Bob who can safely output Y:=YIY:=Y_{I}. In case (1) above, Alice just sends I=1I=1. In case (2), Alice needs to send the entire XX to Bob which requires about the entropy of XX many bits.

Most of the works mentioned above were interested in minimizing the expected length of prefix-free encodings of simulators. This note considers a different (but closely related) minimization problem. The goal here is to minimize the “description length” of II in bits. The cost of a simulator II is

c​o​s​t​(I):=⁡E⁡[log⁡I]cost(I):=\mathop{\mathbb{}}{E}[\log I]

where log\log is base two.

The main purpose of this text is to prove the following theorem, which provides both an upper bound and lower bounds on the cost of simulators.

Theorem.

Let P,QP,Q be two distributions over the same finite set. If D(Q||P)<∞D(Q||P)<\infty then there is a PP to QQ simulator II such that

E[logI]≤D(Q||P)+1.45∥Q−P∥1.\displaystyle\mathop{\mathbb{}}{E}[\log I]\leq D(Q||P)+1.45\|Q-P\|_{1}.

On the other hand, for every PP to QQ simulator II,

E[logI]≥12max{D(Q||P),∥Q−P∥1}.\mathop{\mathbb{}}{E}[\log I]\geq\frac{1}{2}\max\{D(Q||P),\|Q-P\|_{1}\}.

In particular, if D(Q||P)=∞D(Q||P)=\infty then there is no PP to QQ simulator.

Harsha, Jain, McAllester and Radhakrishnan [6] showed that for every P,QP,Q such that D(Q||P)<∞D(Q||P)<\infty, there is a simulator II with cost E[logI]≤D(Q||P)+C\mathop{\mathbb{}}{E}[\log I]\leq D(Q||P)+C where D(Q||P)=∑xQ(x)logQ⁡(x)P⁡(x)D(Q||P)=\sum_{x}Q(x)\log\frac{Q(x)}{P(x)} is the Kullback–Leibler divergence and C>0C>0 is some universal constant. The mechanism behind the proof presented here is different from the mechanism in [6]. Their simulator is based on a carefully constructed rejection sampling procedure. This text uses the Poisson-race mechanism described by Maddison [12] and that also appeared in the work of Li and El Gamal [9] mentioned above. The proof mechanism is informative and it may, e.g., allow to control higher moments of log⁡I\log I.

The upper bound in the theorem comprises two terms. The divergence term D(Q||P)D(Q||P) is dominant when the distributions are far away. The total variation term 1.45​‖Q−P‖11.45\|Q-P\|_{1} is dominant when the distributions are close. The total variation term could be as large as +2.9+2.9, but it is known that it can be replaced by +1+1; see [10] (a simple variant of the proof presented here also leads to this +1+1 bound). The theorem above is strictly stronger than all the results mentioned above when P,QP,Q are close (because all of them had some additive constant term). To the best of my knowledge, the two lower bounds were not observed before.

The lower bound has an interesting consequence. An upper bound of the form cost(I)≤CD(Q||P)cost(I)\leq CD(Q||P) is not achievable in general (where C>0C>0 is any universal constant). For example, when PP is an unbiased coin and QQ is a coin with bias ϵ>0\epsilon>0, we have that D(Q||P)≈ϵ2D(Q||P)\approx\epsilon^{2} but ⁡E​log⁡I≳‖Q−P‖1≈ϵ\mathop{\mathbb{}}{E}\log I\gtrsim\|Q-P\|_{1}\approx\epsilon. In this example, for every II, the cost of II is much larger than D(Q||P)D(Q||P).

The authors of [6] showed a lower bound of D(Q||P)D(Q||P) on the cost of a prefix-free encoding of II. This means that when D(Q||P)D(Q||P) is large, the lower bound stated above is essentially off by a factor of 22. The lower bound E[logI]≥D(Q||P)\mathop{\mathbb{}}{E}[\log I]\geq D(Q||P), however, is not true in general. For example, if PP is a distribution on {0,1}\{0,1\} with P⁡(0)=1−ϵP(0)=1-\epsilon and QQ is defined by Q⁡(0)=1Q(0)=1 then D(Q||P)=ϵln⁡2+O(ϵ2)D(Q||P)=\frac{\epsilon}{\ln 2}+O(\epsilon^{2}) and if II is the first occurrence of 00 then II is a simulator and ⁡E⁡[log⁡I]=ϵ+O⁡(ϵ2)\mathop{\mathbb{}}{E}[\log I]=\epsilon+O(\epsilon^{2}) so that cost(I)≤0.7D(Q||P)cost(I)\leq 0.7D(Q||P).

Here is an intuitive sketch of the mechanism. A Poisson process can be thought of as a cloud of random points on the line ℝ\mathbb{R} (see detailed description and references below). We can (independently) mark each point in the cloud by x∈𝒳x\in\mathcal{X} using PP. So now we have a cloud of marked points. Each mark x∈𝒳x\in\mathcal{X} has density P⁡(x)P(x) in the cloud. This marked cloud represents the distribution PP. Our goal is to shift from PP to QQ. The main point is that Poisson processes are determined by their local densities. Changing densities is straitforward. For example, if we dilate all points marked by xx by a factor of two then we halved their density. More interestingly, if we dilate all points marked xx by P⁡(x)Q⁡(x)\frac{P(x)}{Q(x)}, then their density becomes

P⁡(x)⋅Q⁡(x)P⁡(x)=Q⁡(x).P(x)\cdot\frac{Q(x)}{P(x)}=Q(x).

The new dilated cloud represents QQ. We thus shifted from PP to QQ via a dilation.

Acknowledgement

This note was written following the summer school A-PIC 2026: Probability, Information, Computing held at the African Institute for Mathematical Science in Kigali, Rwanda. I wish to thank Jan Hązła, as well as the participants of the school and the AIMS staff, for their hospitality. I also wish to thank A. El Gamal and C.T. Li for helpful comments. A preliminary draft of this note was written as a preparation for the school with the assistance of ChatGPT.

2. Poisson processes

Here is a very brief introduction to Poisson processes and Poisson races; see for example [2, 8, 7]. A rate-λ\lambda Poisson process on ℝ+\mathbb{R}_{+}

0<T1<T2<…0<T_{1}<T_{2}<\ldots

is characterized by the following properties. First, for every measurable M⊂(0,∞)M\subset(0,\infty) of finite measure |M|<∞|M|<\infty, the count

N⁡(M)=|{i∈ℕ:Ti∈M}|N(M)=\big|\{i\in\mathbb{N}:T_{i}\in M\}\big|

is distributed as

N⁡(M)∼P​o​i​s​(λ​|M|).N(M)\sim Pois(\lambda|M|).

Second, if M1,…,MmM_{1},\ldots,M_{m} are disjoint then the counts N⁡(M1),…,N⁡(Mm)N(M_{1}),\ldots,N(M_{m}) are independent.

We mention a couple of basic properties. If we multiply every TiT_{i} by some fixed α>0\alpha>0 then we get a Poisson process with rate λα\frac{\lambda}{\alpha}. The union of (finitely many) independent Poisson processes is a Poisson process with rate being the sum of the rates.

A key idea for simulation is to use Poisson marking. Assume PP is a distribution on a finite domain 𝒳\mathcal{X} and let X1,X2,…X_{1},X_{2},\ldots be i.i.d. samples from PP. Let 0<T1<T2<…0<T_{1}<T_{2}<\ldots be a rate-11 Poisson process (chosen independently from X1,X2,…X_{1},X_{2},\ldots). For each x∈𝒳x\in\mathcal{X}, we can consider the marked subprocess T1x<T2x<…T^{x}_{1}<T^{x}_{2}<\ldots that is obtained by keeping only the ii’s such that Xi=xX_{i}=x. This marked process is a Poisson process with rate P⁡(x)P(x). In addition, these |𝒳||\mathcal{X}| processes are independent (if P⁡(x)=0P(x)=0, the corresponding process is empty).

3. The simulator

3.1. A Poisson race

The following lemma summarizes the simulator and its correctness.

Lemma.

Let P,QP,Q be two distributions over a finite set 𝒳\mathcal{X} such that D(Q||P)<∞D(Q||P)<\infty. Let X1,X2,…X_{1},X_{2},\ldots be i.i.d. from PP. Let T1<T2<…T_{1}<T_{2}<\ldots be a rate-11 Poisson process on (0,∞)(0,\infty) that is independent of X1,X2,…X_{1},X_{2},\ldots. For each ii, define

Si=TiR⁡(Xi)S_{i}=\frac{T_{i}}{R(X_{i})}

where

R⁡(x):=Q⁡(x)P⁡(x)>0R(x):=\frac{Q(x)}{P(x)}>0

(when Q⁡(Xi)=0Q(X_{i})=0, define Si=∞S_{i}=\infty). Let

I=𝖺𝗋𝗀𝗆𝗂𝗇​Si.I=\mathsf{argmin}\ S_{i}.

Then,

XI∼Q.X_{I}\sim Q.
Proof.

We have |𝒳||{\mathcal{X}}| independent marked Poisson processes (Tix)(T^{x}_{i}). Because D(Q||P)<∞D(Q||P)<\infty, if Q⁡(x)>0Q(x)>0 then P⁡(x)>0P(x)>0. If Q⁡(x)=0Q(x)=0 and P⁡(x)>0P(x)>0 then Six=∞S^{x}_{i}=\infty for all ii. If P⁡(x)=0P(x)=0 then xx never appears. So we can ignore the xx’s such that Q⁡(x)​P​(x)=0Q(x)P(x)=0; they are either fixed or never appear. For a remaining xx, the process (Tix)(T^{x}_{i}) has rate P⁡(x)P(x) so SixS^{x}_{i} is a Poisson process with rate P⁡(x)​R​(x)=Q⁡(x)P(x)R(x)=Q(x).

The processes (Six)(S^{x}_{i}) are independent. Their union is a marked Poisson process with rate 1=∑xQ⁡(x)1=\sum_{x}Q(x). The union process is marked by QQ. Order the elements of the new process as S(1)<S(2)<…S_{(1)}<S_{(2)}<\ldots then the mark of S(1)S_{(1)} is distributed according to QQ. The marks of S(1)S_{(1)} and TIT_{I} are the same by definition so

XI∼Q.X_{I}\sim Q.

The winner II is well-defined because we need to consider only finitely many of the XiX_{i}’s (only the first one of each type x∈𝒳x\in\mathcal{X}). ∎

3.2. Upper bound on cost

It remains to bound ⁡E​log⁡I\mathop{\mathbb{}}{E}\log I from above. First, consider the regular conditional law given {XI=x,SI=s}\{X_{I}=x,S_{I}=s\}. The conditional {SI=s}\{S_{I}=s\} means that S(2)<S(3)<…S_{(2)}<S_{(3)}<\ldots is a rate 1 Poisson process on (s,∞)(s,\infty). The conditional {XI=x}\{X_{I}=x\} means that S(1)S_{(1)} is marked by xx and the rest of the marks are independent of that choice.

In particular, for each y≠xy\neq x, the process (Siy)(S^{y}_{i}) conditioned on {XI=x,SI=s}\{X_{I}=x,S_{I}=s\} is a Poisson process on (s,∞)(s,\infty) with rate Q⁡(y)Q(y). This means that (Tiy)=(R⁡(y)​Siy)(T_{i}^{y})=(R(y)S_{i}^{y}) conditioned on {XI=x,SI=s}\{X_{I}=x,S_{I}=s\} is a Poisson process with rate P⁡(y)P(y) on (s​R​(y),∞)(sR(y),\infty). Conditioned on {XI=x,SI=s}\{X_{I}=x,S_{I}=s\}, the expected number of points of (Tiy)(T^{y}_{i}) before TI=s​R​(x)T_{I}=sR(x) is therefore s​(R⁡(x)−R⁡(y))+​P​(y)s(R(x)-R(y))_{+}P(y).

Summing over yy,

E[I−1|XI=x,SI=s]=∑ysP(y)(R(x)−R(y))+.\mathop{\mathbb{}}{E}[I-1|X_{I}=x,S_{I}=s]=\sum_{y}sP(y)(R(x)-R(y))_{+}.

Because the rate of the process defined by the SiS_{i}’s is 11 and it is independent of the markings,

⁡E⁡[I|XI=x]=⁡Es​[1+∑ys​P​(y)​(R⁡(x)−R⁡(y))+]=1+∑yP⁡(y)​(R⁡(x)−R⁡(y))+.\mathop{\mathbb{}}{E}[I|X_{I}=x]=\mathop{\mathbb{}}{E}_{s}\Big[1+\sum_{y}sP(y)(R(x)-R(y))_{+}\Big]=1+\sum_{y}P(y)(R(x)-R(y))_{+}.

Because

(∑yP⁡(y)​(R⁡(x)−R⁡(y))+)−(∑yP⁡(y)​(R⁡(y)−R⁡(x))+)\displaystyle\Big(\sum_{y}P(y)(R(x)-R(y))_{+}\Big)-\Big(\sum_{y}P(y)(R(y)-R(x))_{+}\Big)
=∑yP⁡(y)​(R⁡(x)−R⁡(y))\displaystyle=\sum_{y}P(y)(R(x)-R(y))
=R⁡(x)−∑yP⁡(y)​Q⁡(y)P⁡(y)=R⁡(x)−1,\displaystyle=R(x)-\sum_{y}P(y)\frac{Q(y)}{P(y)}=R(x)-1,

we can write

1+∑yP⁡(y)​(R⁡(x)−R⁡(y))+\displaystyle 1+\sum_{y}P(y)(R(x)-R(y))_{+} =R⁡(x)+(∑yP⁡(y)​(R⁡(y)−R⁡(x))+).\displaystyle=R(x)+\Big(\sum_{y}P(y)(R(y)-R(x))_{+}\Big).
Remark.

Because XI∼QX_{I}\sim Q, this allows to control the expected value of II as

⁡E⁡[I]=∑xQ⁡(x)​(R⁡(x)+∑yP⁡(y)​(R⁡(y)−R⁡(x))+)≤1+∑xQ2​(x)P⁡(x).\mathop{\mathbb{}}{E}[I]=\sum_{x}Q(x)\Big(R(x)+\sum_{y}P(y)(R(y)-R(x))_{+}\Big)\leq 1+\sum_{x}\frac{Q^{2}(x)}{P(x)}.

Averaging over XI∼QX_{I}\sim Q as well, because log⁡ξ\log\xi is concave,

⁡E⁡[log⁡I]\displaystyle\mathop{\mathbb{}}{E}[\log I] =⁡Ex∼Q​⁡E⁡[log⁡I|x]≤⁡Ex​log⁡(⁡E⁡[I|x]).\displaystyle=\mathop{\mathbb{}}{E}_{x\sim Q}\mathop{\mathbb{}}{E}[\log I|x]\leq\mathop{\mathbb{}}{E}_{x}\log(\mathop{\mathbb{}}{E}[I|x]).

Substituting the above,

⁡E⁡[log⁡I]\displaystyle\mathop{\mathbb{}}{E}[\log I] ≤⁡Ex​log⁡(R⁡(x)+(∑yP⁡(y)​(R⁡(y)−R⁡(x))+))\displaystyle\leq\mathop{\mathbb{}}{E}_{x}\log\Big(R(x)+\Big(\sum_{y}P(y)(R(y)-R(x))_{+}\Big)\Big)
=D(Q||P)+(∑xQ(x)log(1+(∑yP(y)(R⁡(y)−R⁡(x))+R⁡(x))).\displaystyle=D(Q||P)+\Big(\sum_{x}Q(x)\log\Big(1+\Big(\sum_{y}P(y)\frac{(R(y)-R(x))_{+}}{R(x)}\Big)\Big).

It remains to bound the second term. Because log2⁡(1+ξ)≤ξln⁡2\log_{2}(1+\xi)\leq\frac{\xi}{\ln 2} for ξ≥0\xi\geq 0,

∑xQ⁡(x)​log⁡(1+(∑yP⁡(y)​(R⁡(y)−R⁡(x))+R⁡(x))CLOSE\displaystyle\sum_{x}Q(x)\log\Big(1+\Big(\sum_{y}P(y)\frac{(R(y)-R(x))_{+}}{R(x)}\Big)
≤1ln⁡2​∑x,yQ⁡(x)​P​(y)​(R⁡(y)−R⁡(x))+R⁡(x)\displaystyle\leq\frac{1}{\ln 2}\sum_{x,y}Q(x)P(y)\frac{(R(y)-R(x))_{+}}{R(x)}
=1ln⁡2​∑x,yP⁡(x)​P​(y)​(R⁡(y)−R⁡(x))+.\displaystyle=\frac{1}{\ln 2}\sum_{x,y}P(x)P(y)(R(y)-R(x))_{+}.

By “symmetry” between x,yx,y,

∑x,yP⁡(x)​P​(y)​(R⁡(y)−R⁡(x))+\displaystyle\sum_{x,y}P(x)P(y)(R(y)-R(x))_{+} =12​∑x,yP⁡(x)​P​(y)​|R⁡(y)−R⁡(x)|\displaystyle=\frac{1}{2}\sum_{x,y}P(x)P(y)|R(y)-R(x)|
=12​∑x,yP⁡(x)​P​(y)​|R⁡(y)−1+1−R⁡(x)|\displaystyle=\frac{1}{2}\sum_{x,y}P(x)P(y)|R(y)-1+1-R(x)|
≤∑x,yP⁡(x)​P​(y)​|R⁡(x)−1|=‖Q−P‖1.\displaystyle\leq\sum_{x,y}P(x)P(y)|R(x)-1|=\|Q-P\|_{1}.

Altogether,

E[logI]≤D(Q||P)+‖Q−P‖1ln⁡2\displaystyle\mathop{\mathbb{}}{E}[\log I]\leq D(Q||P)+\frac{\|Q-P\|_{1}}{\ln 2}

4. Lower bounds

Assume X1,X2,…X_{1},X_{2},\ldots are i.i.d. from PP and that II is such that XI∼QX_{I}\sim Q.

4.1. Total variation bound

For each xx,

Pr[XI=x,I=1]≤Pr[X1=x]=P(x)\Pr[X_{I}=x,I=1]\leq\Pr[X_{1}=x]=P(x)

and

Pr[XI=x,I=1]≤Pr[XI=x]=Q(x).\Pr[X_{I}=x,I=1]\leq\Pr[X_{I}=x]=Q(x).

So,

Pr[I=1]≤∑xmin{P(x),Q(x)}=1−‖Q−P‖12.\Pr[I=1]\leq\sum_{x}\min\{P(x),Q(x)\}=1-\frac{\|Q-P\|_{1}}{2}.

If I>1I>1 then log⁡I≥1\log I\geq 1 so

E[logI]≥Pr[I>1]≥‖Q−P‖12.\displaystyle\mathop{\mathbb{}}{E}[\log I]\geq\Pr[I>1]\geq\frac{\|Q-P\|_{1}}{2}.

4.2. Divergence bound

For every xx such that Q⁡(x)>0Q(x)>0,

Q(x)Pr[I=i|XI=x]=Pr[XI=x,I=i]≤Pr[Xi=x]≤P(x)Q(x)\Pr[I=i|X_{I}=x]=\Pr[X_{I}=x,I=i]\leq\Pr[X_{i}=x]\leq P(x)

or

Pr⁡[I=i|XI=x]≤1R⁡(x)\Pr[I=i|X_{I}=x]\leq\frac{1}{R(x)}

where R⁡(x)=Q⁡(x)P⁡(x)R(x)=\frac{Q(x)}{P(x)}.

If R⁡(x)≤1R(x)\leq 1 then

⁡E⁡[log⁡I|XI=x]≥0≥12​log⁡R⁡(x).\mathop{\mathbb{}}{E}[\log I|X_{I}=x]\geq 0\geq\frac{1}{2}\log R(x).

If R⁡(x)>1R(x)>1 then argue as follows. Out of all distributions of YY on {1,2,…}\{1,2,\ldots\} such that Pr[Y=y]≤1R\Pr[Y=y]\leq\frac{1}{R} for all yy, the one that minimizes ⁡E​log⁡Y\mathop{\mathbb{}}{E}\log Y gives mass 1R\frac{1}{R} for the first M:=⌊R⌋≥1M:=\lfloor R\rfloor\geq 1 elements and the rest of the mass 1−MR≤1R1-\frac{M}{R}\leq\frac{1}{R} is given to the element M+1M+1. For this distribution,

⁡E⁡[log⁡Y]\displaystyle\mathop{\mathbb{}}{E}[\log Y] =(∑i≤M1R​log⁡i)+(1−MR)​log⁡(M+1)\displaystyle=\Big(\sum_{i\leq M}\frac{1}{R}\log i\Big)+\Big(1-\frac{M}{R}\Big)\log(M+1)
=log⁡(M!)+(R−M)​log⁡(M+1)R.\displaystyle=\frac{\log(M!)+(R-M)\log(M+1)}{R}.

Write R=M+θR=M+\theta with 0≤θ<10\leq\theta<1. For an integer nn, we have log⁡(n!)≥n2​log⁡n\log(n!)\geq\frac{n}{2}\log n because k⁡(n+1−k)≥nk(n+1-k)\geq n. So,

log⁡(M!)+(R−M)​log⁡(M+1)\displaystyle\log(M!)+(R-M)\log(M+1) =log⁡(M!)+θ​log⁡(M+1)\displaystyle=\log(M!)+\theta\log(M+1)
=(1−θ)​log⁡(M!)+θ​log⁡((M+1)!)\displaystyle=(1-\theta)\log(M!)+\theta\log((M+1)!)
≥(1−θ)​M2​log⁡(M)+θ​M+12​log⁡(M+1).\displaystyle\geq(1-\theta)\frac{M}{2}\log(M)+\theta\frac{M+1}{2}\log(M+1).

Because ξ​log⁡ξ\xi\log\xi is convex,

log⁡(M!)+(R−M)​log⁡(M+1)\displaystyle\log(M!)+(R-M)\log(M+1)
≥12​((1−θ)​M+θ⁡(M+1))​log⁡((1−θ)​M+θ⁡(M+1))=12​R​log⁡R.\displaystyle\geq\frac{1}{2}\big((1-\theta)M+\theta(M+1)\big)\log\big((1-\theta)M+\theta(M+1)\big)=\frac{1}{2}R\log R.

We can conclude that also if R⁡(x)>1R(x)>1,

⁡E⁡[log⁡I|XI=x]≥12​log⁡R⁡(x).\mathop{\mathbb{}}{E}[\log I|X_{I}=x]\geq\frac{1}{2}\log R(x).

Taking expectation,

⁡E⁡[log⁡I]=⁡Ex∼Q​⁡E⁡[log⁡I|x]≥12​⁡E​log⁡R⁡(x)=D(Q||P)2.\mathop{\mathbb{}}{E}[\log I]=\mathop{\mathbb{}}{E}_{x\sim Q}\mathop{\mathbb{}}{E}[\log I|x]\geq\frac{1}{2}\mathop{\mathbb{}}{E}\log R(x)=\frac{D(Q||P)}{2}.

References

  • [1] A. Block and Y. Polyanskiy (2023) The sample complexity of approximate rejection sampling with applications to smoothed online learning. In The Thirty Sixth Annual Conference on Learning Theory, pp. 228–273. Cited by: §1.
  • [2] D. J. Daley and D. Vere-Jones (2003) An introduction to the theory of point processes, volume i: elementary theory and methods. 2 edition, Probability and Its Applications, Springer, New York, NY. External Links: Document, ISBN 9780387955414 Cited by: §2.
  • [3] A. El Gamal and Y. Kim (2011) Network information theory. Cambridge university press Cambridge, UK. Cited by: §1.
  • [4] D. Goc and G. Flamich (2024) On channel simulation with causal rejection samplers. In 2024 IEEE International Symposium on Information Theory (ISIT), pp. 1682–1687. Cited by: §1.
  • [5] B. Hajek and M. Pursley (1979) Evaluation of an achievable rate region for the broadcast channel. IEEE Transactions on Information Theory 25 (1), pp. 36–46. Cited by: §1.
  • [6] P. Harsha, R. Jain, D. A. McAllester, and J. Radhakrishnan (2007) The communication complexity of correlation. In CCC 2007, pp. 10–23. Cited by: §1, §1, §1.
  • [7] J. F. C. Kingman (1993) Poisson processes. Oxford Studies in Probability, Vol. 3, Clarendon Press, Oxford. External Links: ISBN 9780198536932 Cited by: §2.
  • [8] G. Last and M. Penrose (2017) Lectures on the poisson process. Cambridge University Press. External Links: Document, ISBN 9781107458437 Cited by: §2.
  • [9] C. T. Li and A. E. Gamal (2018) Strong functional representation lemma and applications to coding theorems. IEEE Transactions on Information Theory 64 (11), pp. 6967–6978. Cited by: §1, §1.
  • [10] C. T. Li (2024) Pointwise redundancy in one-shot lossy compression via poisson functional representation. arXiv:2401.14805. Cited by: §1.
  • [11] J. Liu and S. Verdu (2018) Rejection sampling and noncausal sampling under moment constraints. In 2018 IEEE International Symposium on Information Theory (ISIT), pp. 1565–1569. Cited by: §1.
  • [12] C. J. Maddison (2016) A poisson process model for monte carlo. In Perturbations, Optimization, and Statistics, pp. 193–232. Cited by: §1.
  • [13] J. von Neumann (1951) Various techniques used in connection with random digits. In Monte Carlo Method, National Bureau of Standards Applied Mathematics Series, pp. 36–38. Cited by: §1.
  • [14] A. D. Wyner (1975) The common information of two dependent random variables. IEEE Transactions on Information Theory 21 (2), pp. 163–179. Cited by: §1.